a wooden fence is to be built around a 28 by 78 lot how many meters for the fence will be needed ? if the wood for the fence costs 36.50 per meter what will the wood for the fence cost?
Answer by macston(5194)

Perimeter=2(L+W)=2(28m+78m)=2(106m)=212m ANSWER 1: It will take 212 meters of fence.
(212 meters)(36.50/meter)= 7738 ANSWER 2: The fence will cost 7738 (currency units not given).